# EvalMuse-40K: A Reliable and Fine-Grained Benchmark with Comprehensive Human Annotations for Text-to-Image Generation Model Evaluation

### v1

@@ 3.4 Data Statistics and Reliability Analysis
To summarize, we generate 40K images using 4K prompts based on 20 T2I models. For each image-text pair, we perform multiple annotations, including image-text alignment scores, element matching, and structural issues, achieving nearly 1M annotations. We employ novel sampling and generation strategies to ensure a balanced and diverse set of prompts. Additionally, we implement multi-round annotation and revision processes to guarantee the reliability of the annotated data.
@@
To analyze the agreement of the scoring data among annotators, we calculate the maximum score difference for each image-text pair and plot it as a histogram in Fig. 2(b). It can be seen that 75% of the samples show a score difference of less than 1 point. It is worth noting that for score differences of 2 or more, we obtain double annotations (from 3 to 6) by re-annotating, further reducing the inter-annotator disagreement.
For fine-grained annotation, we perform statistics on the quantity and alignment scores of elements based on their respective categories. It can be observed from Fig. 2(c) that the overall alignment scores for most categories are around 50%, ensuring a balanced distribution of positive and negative samples in fine-grained annotation. Additionally, we found that the images generated by AIGC models exhibited relatively poor consistency with the text in terms of counting, spatial aspects, and activity aspects.
@@ 4 Methods for Alignment Evaluation
In this section, we introduce two methods for evaluating image-text alignment in AIGC tasks. The proposed FGA-BLIP2 can achieve Fine-Grained Alignment evaluation through end-to-end training. The proposed PN-VQA can perform the zero-shot evaluation using a Positive-Negative VQA manner.
@@ 4.1 End-to-End Scoring Model: FGA-BLIP2
Most current scoring models for image-text alignment tasks are trained using reward model objectives, which can learn human preferences between different images generated for the same prompt. Since alignment scores are not available during training, these methods do not always produce values that accurately reflect the degree of image-text alignment. Benefiting from the large number of image-text alignment scores labeled in EvalMuse-40K, we can train the mapping between image-text pairs and scores end-to-end.
Additionally, FGA-BLIP2 achieves fine-grained alignment evaluation by jointly training the overall and elemental alignment scores (see Fig. 3). We adopt the setting of Image-Text Matching (ITM) in BLIP2 [23], where the query and text are concatenated and then processed through cross-attention with the image. This yields query embedding and text embedding outputs. The final alignment score is obtained by a two-class linear classifier, where the query part is averaged to produce the overall alignment score, while the text part at each corresponding position provides the element-specific alignment scores. Since not all tokens in the text are highly relevant to the alignment task, we introduce an additional operation to predict valid tokens. The text is first inputted into a self-attention layer to extract text features, which then are passed to an MLP to predict a mask that represents the validity of each text token.
Meanwhile, we observed that some prompts in the dataset are either too simple or overly complex, resulting in minimal differences in alignment scores across images generated by different models. Such data may lead the model to focus more on prompt complexity than on the actual alignment level of the image-text pairs during training. We therefore design a variance-weighted optimization strategy for image-text alignment task. We calculate the variance $\sigma(p)$ of the alignment scores across different images generated from the same prompt and use this to adjust the loss weights of different prompts during training.
@@
where weighting parameters are set to $\lambda=0.1$ and $\eta=0.1$. $L_{os}$ denotes the L1 loss between the predicted overall alignment score and human annotation. $L_{es}$ denotes the L1 loss between the predicted element score and human fine-grained annotation. $L_{mask}$ denotes the L1 loss between the predicted valid text token and the real elements. When using variance to weight the loss function, a larger $\sigma(p)$ for an image-text pair results in a higher loss. This approach encourages the model to focus more on samples with greater differences in image-text matching scores under the same prompt, helping the model better understand and evaluate alignment level between images and text.
Figure 3: Structure of FGA-BLIP2. $L_{\text{os}}$ represents the overall alignment score loss, $L_{\text{es}}$ is the element alignment score loss, and $L_{\text{mask}}$ is the loss for predicting valid elements. FGA-BLIP2 jointly optimizes these losses to achieve fine-grained evaluation.
@@ 4.2 Fine-Grained Zero-shot Alignment: PN-VQA
Given the good performance of MLLMs, we follow the previous methods [16, 45] of using question answering for fine-grained evaluation. In comparison to the previous methods, we only generate simple yes/no judgment questions to evaluate whether the elements in the text are accurately represented in the generated images. By using straightforward yes/no questions, we effectively capture the VQA model’s probability of outputting correct or incorrect answers.
@@
We then average to get the fine-grained alignment score as $S_{e}=(P_{T}+1-P_{F})/2$. This positive-negative questioning method enables MLLMs to perform a more robust fine-grained evaluation of generated images. Additionally, we found that incorporating the original prompt into the question allows MLLMs to better assess the alignment of elements with the image.
@@ 5.1 Experimental Setup
We use one-quarter of the samples from EvalMuse-40K as the test set, ensuring no overlap in prompts between the training and test sets. The test set includes 500 real prompts and 500 synthetic prompts. We train FGA-BLIP2 on the training set and then test existing evaluation methods along with our FGA-BLIP2 and PN-VQA on the test set. Additionally, we use GenAI-Bench [20], TIFA [16], and RichHF [25] datasets to validate the generalization capability of our fine-tuned model FGA-BLIP2.
Training Settings. We use the BLIP-2 [23] model fine-tuned on the COCO [27] for the retrieval task as initialization weights. We follow BLIP-2’s training setup to set the maximum learning rate to $1e^{-5}$ and the minimum learning rate to 00 and train for 5 epochs on an A100 GPU.
@@ 5.2 Results on Automated Metrics
5.2.1 Quantitative Analysis
We evaluate our proposed metrics from two aspects: overall alignment evaluation and fine-grained evaluation.
Overall Alingment. In Tab. 2, we report the results of FGA-BLIP2 and other state-of-the-art methods on multiple benchmarks. The methods used here can directly output alignment scores from image-text pairs without extra operations like VQA. It is evident that FGA-BLIP2 performs excellently not only on the EvalMuse-40K test set but also exhibits higher correlation with human annotations on GenAI-Bench [20], TIFA800 [16], and the alignment score sections of RichHF [25] compared to other methods.
@@ 5.2.1 Quantitative Analysis
For different elemental skills in fine-graining such as counting and color, we also compare multiple fine-graining evaluation methods, and both FGA-BLIP2 and PN-VQA achieve good performance. Meanwhile, we select 200 representative prompts and use our FGA-BLIP2 to evaluate the alignment of image-text pairs generated by over twenty T2I models. We found that current proprietary models such as Midjourney [15] and Dreamina [9] generally performed better. Additionally, we evaluate MLLMs’ ability to detect structural problems using EvalMuse-40K’s structural labels. Due to space constraints, the comparison results for all element skills, the ranking of T2I models’ alignment ability by FGA-BLIP2, and the exploration of different MLLMs’ structural problem detection abilities are provided in the supplementary material.
5.2.2 Qualitative Analysis
In Fig. 4, we present the prediction scores of FGA-BLIP2 and VQAScore [21] across multiple image-text pairs. It can be seen that FGA-BLIP2 produces scores that are more consistent with human annotations. For fine-grained evaluation, as shown in Fig. 5, our proposed methods are able to more accurately determine whether the elements in the prompt are represented in the generated image.
5.2.3 Ablation Study
The ablation study for FGA-BLIP2 is shown in Tab. 2. When evaluating using the overall alignment scores (os) output by FGA-BLIP2, our model achieves better correlation across four datasets compared to the method without the variance optimization strategy (w/o var). We also report the fine-grained element scores (es) output by FGA-BLIP2 and take the average (avg) as the image-text alignment score for evaluating the model’s performance on the datasets. Combining and weighting os and es_avg as the final score leads to improved performance on GenAI-Bench and TIFA.
@@ 7 Overview
In this supplementary material, we provide data collection details, additional experiment results, and T2I model alignment evaluation. In data collection details, we detail the classification and sampling of real user prompts in Sections 8.1 and 8.2, ensuring diversity and balance among the real prompts. We then describe the specific method for generating synthetic prompts in Section 8.3 and detail the element splitting and question generation processes in Sections 8.4 and  8.5. Regarding additional experimental results, we report the performance of FGA-BLIP2 and PN-VQA on fine-grained specific skills (see Tab. 6) and provide additional visualizations in Sections 9.1 and  9.2. We evaluate the structure-aware capabilities of MLLMs using the structural labels in EvalMuse-40K in Section 9.3. Finally, in T2I model alignment evaluation, we employ FGA-BLIP2 to assess more than twenty widely used T2I models. The evaluation processes and results are presented in Section 10 and Tab. 10.
@@ 8.1 Real Prompt Categorization
Specifically, subject categories are divided into the following types: Artifacts, World Knowledge, People, Outdoor Scenes, Illustrations, Vehicles, Food & Beverage, Arts, Abstract, Produce & Plants, Indoor Scenes, Animals, and Idioms. Logical relationships are divided into Position Relationship, Number, Color, Writing & Symbols, Perspective, and Anti-reality. Image style dimension includes Material, Genre, Design, Photography & Cinema, and Artist & Works. BERT embeddings are used to represent the semantic information of the prompts.
The prompt collection process is as follows. We first randomly select 100K prompts from DiffusionDB as the source dataset. GPT-4 [1] is then utilized to label the three dimensions of subject category, logical relationship, and image style of the prompt representation, meanwhile allowing the prompt to have multiple labels in a given dimension. In Tab. 5, we show examples of real prompts from various categories. The semantic dimensions represented by the prompt’s Bert embedding are then categorized into seven types using the K-Means clustering algorithm.
@@ 8.2 Prompt Sample Strategy
which means to select $N$ samples from the original data such that their distribution closely resembles $D$. The above minimization can be solved by using a set of auxiliary vectors $\mathbf{Z}=\{\mathbf{z}_{m}\}_{m=1}^{M}$ with $\mathbf{z}_{m}\in\mathbb{R}^{H_{m}}_{+}$, in order to handle the absolute values of the $L_{1}$ norm [39]:
@@
Figure 6: Distribution of data before and after sampling in four dimensions using MILP strategy. It can be observed that through MILP sampling, the distribution of prompts across various dimensions is more balanced.
@@ 8.3 Sythetic Data Generation
We generate 2K prompts by using keywords combination and GPT-4. These prompts can be categorized into six different classes.
@@
Naive Object Count. We prepare a table that contains hundreds of object names. These objects can be classified into animal, human, object-household, object-electronic, object-cloth, object-others, object-vehicle, and food. To generate the prompt, we randomly select two numbers from 0-6 and two objects, then concatenate them together. For example, given sampled number 1 and some, objects cat and dogs, the generated prompt should be “1 cat and some dogs”.
@@
Object Color Material. We randomly select two objects and two colors from our keyword table. We use GPT-4 to generate a prompt following the fixed format which contains color, object, and material. The prompt is “Generate a natural and reasonable prompt following the format [color] [adj-material][noun] and [color][adj-material][noun], given noun {xx} and color {xx}.” For example, given color red and green and the object cat and dog, the generated prompt can be “a furry red cat and a green plastic dog”
@@
Object Location Weather Timeline Style. We randomly select one object name, one location keyword and one style keyword from our keyword table. The prompt is generated by GPT-4 by giving the following prompt: Generate a prompt following the format A/an [noun][adp] [location], [style], [weather], [time], given noun {xx}, location {xx} and style {xx}. For example, given noun crew, location river, style 2D, the generated prompt should be “A crew at the River, 2D style, sunny, afternoon”.
@@
Attribute Object Activity Perspective View. We randomly select one object name from the human and animal class, then use GPT-4 to generate a verb and a noun that correctly match the selected object. The prompt is shown as “Generate a natural and reasonable prompt following the format [nounA] [verb][nounB], given nounA {xx}.” For example, given nounA: miner, the generated prompt should be “miner digs ore”. After that, we add one perspective view-related keyword to the end of the phrase. The keyword is randomly selected from our keyword lists. The final generated prompt should follow the format “miner digs ore, panorama”
@@
Naive Writing Symbols. We collect a set of words and symbols. We randomly select one element from the words-symbol set and select one word style keyword from [“Regular script font”, “Times New Roman font”, “Art style”, “Landscape orientation”, “Portrait orientation”]. Then we set the prompt as “a blackboard with text {xx} on it, {xx} style”. For example, given the word “love” and style keyword “Art style”, the prompt should be “a blackboard with text ’love’ on it, Art Style”
@@
Attribute Spatial Composition. We randomly select two object names, one spatial keyword, and one attribute keyword, then use GPT-4 to generate the prompt that reflects the spatial and attribute relationship between these objects. The prompt is shown as “Generate a natural and reasonable prompt following the format [nounA] [position] [noun B] and the [noun A] [adj] the [noun B], given noun {xx}. [position] denotes the positional relationship between objects and [adj] denotes the contrasting relationship between objects.” For example, given the noun boy and girl, the generated prompt may be: “a boy is beside a girl, and the boy is taller than the girl.”

### v2

@@ 3.4 Data Statistics and Reliability Analysis
We generate 40K images using 4K prompts based on 20 T2I models. For each image-text pair, we perform multiple annotations, including image-text alignment scores, element matching, and structural issues, achieving nearly 1M annotations. We employ novel sampling and generation strategies to ensure a balanced and diverse set of prompts. Additionally, we implement multi-round annotation and revision processes to guarantee the reliability of the annotated data.
@@
To analyze the agreement of the scoring data among annotators, we calculate the maximum score difference for each image-text pair and plot it as a histogram in Fig. 2(b). It can be seen that 75% samples show a score difference of less than 1 point. It is worth noting that for score differences of 2 or more, we obtain double annotations (from 3 to 6) by re-annotating, further reducing the inter-annotator disagreement.
For fine-grained annotation, we perform statistics on quantity and alignment scores of elements based on their respective categories. It can be observed from Fig. 2(c) that the overall alignment scores for most categories are around 50%, ensuring a balanced distribution of positive and negative samples in fine-grained annotation. Additionally, we observe that the images generated by AIGC models exhibit relatively poor consistency with the text in aspects of counting, shape, and activity.
@@ 4 Methods for Alignment Evaluation
In this section, we introduce two methods for evaluating image-text alignment in AIGC tasks. The proposed FGA-BLIP2 can achieve Fine-Grained Alignment evaluation through end-to-end training. The proposed PN-VQA can perform zero-shot evaluation using a Positive-Negative VQA manner.
@@ 4.1 End-to-End Scoring Model: FGA-BLIP2
Most scoring models for image-text alignment tasks are trained using reward model objectives, which learn human preferences between different images generated for the same prompt. Since alignment scores are not available during training, these methods often fail to produce values that accurately reflect the degree of image-text alignment. Benefiting from the extensive image-text alignment scores labeled in EvalMuse-40K, we can train an end-to-end mapping between image-text pairs and their corresponding scores.
Additionally, FGA-BLIP2 achieves fine-grained alignment evaluation by jointly training the overall and elemental alignment scores (see Fig. 3). We adopt the setting of Image-Text Matching (ITM) in BLIP2 [23], where the query and text after embedding are concatenated and then processed through cross-attention with the image. This yields query embedding and text embedding outputs. The final alignment score is obtained by a two-class linear classifier, where the query part is averaged to produce the overall alignment score, while the text part at each corresponding position provides the element-specific alignment scores. Since not all tokens in the text are highly relevant to the alignment task, we introduce an additional operation to predict valid tokens. The text is first fed to a self-attention layer to extract text features, which then are passed to an MLP to predict a mask that represents the validity of each text token.
We observe that some prompts in the dataset are either too simple or overly complex, resulting in minimal differences in alignment scores across images generated by different models. Such data may lead the model to focus more on prompt complexity than on the actual alignment level of the image-text pairs during training. We therefore design a variance-weighted optimization strategy for the image-text alignment task. We calculate the variance $\sigma(p)$ of the alignment scores across different images generated using the same prompt and use this to adjust the loss weights of different prompts during training.
@@
where weighting parameters are set to $\lambda=0.1$ and $\eta=0.1$. $L_{os}$ denotes the L1 loss between the predicted overall alignment score and human annotation. $L_{es}$ denotes the L1 loss between the predicted element score and human fine-grained annotation. $L_{mask}$ denotes the L1 loss between the predicted valid text token and the real elements. When using variance to weight the loss function, a larger $\sigma(p)$ for an image-text pair results in a higher loss. This approach encourages the model to focus more on samples with greater differences in image-text matching scores under the same prompt, helping the model better understand and evaluate the alignment level between image and text.
Figure 3: Structure of FGA-BLIP2. $m$ and $n$ denote query and text token lengths, respectively; $s_{o}$ and $s_{e}$ represent overall and element scores, respectively; $L_{\text{os}}$ represents the overall alignment score loss, $L_{\text{es}}$ is the element alignment score loss, and $L_{\text{mask}}$ is the loss for predicting valid elements. FGA-BLIP2 jointly optimizes these losses to achieve fine-grained evaluation.
@@ 4.2 Fine-Grained Zero-Shot Alignment: PN-VQA
Given the good performance of MLLMs, we follow the previous methods [16, 45] of using question answering for fine-grained evaluation. In comparison to the previous methods, we only generate simple yes/no judgment questions to evaluate whether the elements in the text are accurately represented in the generated image. By using straightforward yes/no questions, we effectively capture the VQA model’s probability of outputting correct or incorrect answers.
@@
We then average to obtain the fine-grained alignment score as $S_{e}=(P_{T}+1-P_{F})/2$. This positive-negative questioning method enables MLLMs to perform a more robust fine-grained evaluation of generated images. Additionally, we found that incorporating the original prompt into the question enables MLLMs to more effectively assess the alignment of elements with the image.
@@ 5.1 Experimental Setup
We use one-quarter of the samples from EvalMuse-40K as test set, ensuring no overlap in prompts between the training and test sets. The test set includes 500 real prompts and 500 synthetic prompts. We train FGA-BLIP2 on the training set and then test existing evaluation methods along with our FGA-BLIP2 and PN-VQA on the test set. Additionally, we use GenAI-Bench [20], TIFA [16], and RichHF [25] datasets to validate the generalization capability of our fine-tuned model FGA-BLIP2.
Training Settings. We use the BLIP-2 [23] model fine-tuned on the COCO [27] for the retrieval task as initialization weights. We follow BLIP-2’s training setup to set the maximum learning rate to $1e^{-5}$ and the minimum learning rate to 00 and train 5 epochs on an A100 GPU.
@@
Figure 5: Visualization of fine-grained alignment scores from different methods. In each figure, the question and answer are generated based on the element in the prompt and the score represents the probability of the element in the prompt matching the image (indicates a match, and indicates a mismatch). Here, both VQ2 [45] and PN-VQA use Qwen2-VL [40] for VQA task. “Human” refers to the average fine-grained scores from multiple annotators. It can be seen that our FGA-BLIP2 and PN-VQA yield results consistent with human annotations.
@@ 5.2 Results on Automated Metrics
Quantitative Analysis. We evaluate our proposed metrics from two aspects: overall alignment evaluation and fine-grained evaluation.
Overall Alignment. In Tab. 2, we report the results of FGA-BLIP2 and the state-of-the-art methods on multiple benchmarks. The methods used here can directly output alignment scores from image-text pairs without extra operations like VQA. FGA-BLIP2 performs excellently not only on the EvalMuse-40K test set but also exhibits higher correlation with human annotations on GenAI-Bench [20], TIFA800 [16], and the alignment score sections of RichHF [25] compared to other methods.
@@
For different elemental skills in fine-graining such as counting and color, we also compare multiple fine-graining evaluation methods, and both FGA-BLIP2 and PN-VQA achieve good performance. Meanwhile, we select 200 representative prompts and use our FGA-BLIP2 to evaluate the alignment of image-text pairs generated by over 20 T2I models. We found that current proprietary models such as Midjourney [15] and Dreamina [9] generally performed better. Additionally, we evaluate MLLMs’ ability to detect structural problems using EvalMuse-40K’s structural labels. Due to space constraints, the comparison results for all element skills, the ranking of T2I models’ alignment ability by FGA-BLIP2, and the exploration of different MLLMs’ structural problem detection abilities are provided in the supplementary material.
Qualitative Analysis. In Fig. 4, we present the prediction scores of FGA-BLIP2 and VQAScore [21] across multiple image-text pairs. It can be seen that FGA-BLIP2 produces scores that are more consistent with human annotations. For fine-grained evaluation, as shown in Fig. 5, our proposed methods are able to more accurately determine whether the elements in the prompt are represented in the generated image.
5.3 Ablation Study
The ablation study for FGnA-BLIP2 is shown in Tab. 2. When evaluating using the overall alignment scores (os) output by FGA-BLIP2, our model achieves better correlation across four datasets compared to the method without the variance optimization strategy (w/o var). We also report the fine-grained element scores (es) output by FGA-BLIP2 and take the average (avg) as the image-text alignment score for evaluating the model’s performance on the datasets. Combining and weighting os and es_avg as the final score leads to improved performance on GenAI-Bench and TIFA.
@@ 7 Overview
In this supplementary material, we provide data collection details, additional experiment results, and T2I model alignment evaluation. In data collection details, we detail the classification and sampling of real user prompts in Sections 8.1 and 8.2, ensuring diversity and balance among the real prompts. We then describe the specific method for generating synthetic prompts in Section 8.3 and detail the element splitting and question generation processes in Sections 8.4 and  8.5. Regarding additional experimental results, we report the performance of FGA-BLIP2 and PN-VQA on fine-grained specific skills (see Tab. 6) and provide additional visualizations in Sections 9.1 and  9.2. We evaluate the structure-aware capabilities of MLLMs using the structural labels in EvalMuse-40K in Section 9.3. Finally, in T2I model alignment evaluation, we employ FGA-BLIP2 to assess more than 20 widely used T2I models. The evaluation processes and results are presented in Section 10 and Tab. 10.
@@ 8.1 Real Prompt Categorization
Specifically, subject categories are divided into the following types: Artifacts, World Knowledge, People, Outdoor Scenes, Illustrations, Vehicles, Food & Beverage, Arts, Abstract, Produce & Plants, Indoor Scenes, Animals, and Idioms. Logical relationships are divided into Position Relationship, Number, Color, Writing & Symbols, Perspective, and Anti-reality. Image style includes Material, Genre, Design, Photography & Cinema, and Artist & Works. BERT embeddings are used to represent the semantic information of the prompts.
The prompt collection process is as follows. We first randomly select 100K prompts from DiffusionDB as the source dataset. GPT-4 [1] is then utilized to label the three dimensions of subject category, logical relationship, and image style of the prompt representation, meanwhile allowing the prompt to have multiple labels in a given dimension. In Tab. 5, we show examples of real prompts from various categories. The semantic dimensions represented by the prompt’s BERT embeddings are then categorized into seven types using the K-Means clustering algorithm.
@@ 8.2 Prompt Sample Strategy
which involves selecting $N$ samples from the original data such that their distribution closely aligns with $D$. The above minimization can be solved by using a set of auxiliary vectors $\mathbf{Z}=\{\mathbf{z}_{m}\}_{m=1}^{M}$ with $\mathbf{z}_{m}\in\mathbb{R}^{H_{m}}_{+}$, in order to handle the absolute values of the $L_{1}$ norm [39]:
@@
Figure 6: Distribution of data before (top) and after (bottom) sampling in four dimensions using MILP strategy. It can be observed that through MILP sampling, the distribution of prompts across various dimensions is more balanced.
@@ 8.3 Sythetic Data Generation
We generate 2K prompts by using keyword combination and GPT-4. These prompts can be categorized into six different classes.
@@
Naive Object Count. We prepare a table that contains hundreds of object names. These objects can be classified into animal, human, object-household, object-electronic, object-cloth, object-others, object-vehicle, and food. To generate the prompt, we randomly select two numbers from 0-6 and two objects, then concatenate them. For example, given sampled number 1 and some, objects cat and dogs, the generated prompt should be “1 cat and some dogs”.
@@
Object Color Material. We randomly select two objects and two colors from our keyword table. We use GPT-4 to generate a prompt following the fixed format which contains color, object, and material. The prompt is “Generate a natural and reasonable prompt following the format [color] [adj-material][noun] and [color][adj-material][noun], given noun {xx} and color {xx}.” For example, given color red and green and the object cat and dog, the generated prompt can be “a furry red cat and a green plastic dog”.
@@
Object Location Weather Timeline Style. We randomly select one object name, one location keyword, and one style keyword from our keyword table. The prompt is generated by GPT-4 by giving the following prompt: Generate a prompt following the format A/an [noun][adp] [location], [style], [weather], [time], given noun {xx}, location {xx} and style {xx}. For example, given noun crew, location river, style 2D, the generated prompt should be “A crew at the River, 2D style, sunny, afternoon”.
@@
Attribute Object Activity Perspective View. We randomly select one object name from the human and animal class, then use GPT-4 to generate a verb and a noun that correctly match the selected object. The prompt is shown as “Generate a natural and reasonable prompt following the format [nounA] [verb][nounB], given nounA {xx}.” For example, given nounA: miner, the generated prompt should be “miner digs ore”. After that, we add one perspective view-related keyword to the end of the phrase. The keyword is randomly selected from our keyword lists. The final generated prompt should follow the format “miner digs ore, panorama”.
@@
Naive Writing Symbols. We collect a set of words and symbols. We randomly select one element from the words-symbol set and select one word style keyword from [“Regular script font”, “Times New Roman font”, “Art style”, “Landscape orientation”, “Portrait orientation”]. Then we set the prompt as “a blackboard with text {xx} on it, {xx} style”. For example, given the word “love” and style keyword “Art style”, the prompt should be “a blackboard with text ’love’ on it, Art Style”.
@@
Attribute Spatial Composition. We randomly select two object names, one spatial keyword, and one attribute keyword, then use GPT-4 to generate the prompt that reflects the spatial and attribute relationship between these objects. The prompt is shown as “Generate a natural and reasonable prompt following the format [nounA] [position] [noun B] and the [noun A] [adj] the [noun B], given noun {xx}. [position] denotes the positional relationship between objects and [adj] denotes the contrasting relationship between objects.” For example, given the noun boy and girl, the generated prompt may be: “a boy is beside a girl, and the boy is taller than the girl”.